An anti-aircraft shell is fired vertically upward with a muzzle velocity of 100m/s. Calculate the maximum height it can attain

1
Expert's answer
2021-12-16T18:03:42-0500

According to the energy conservation law, the initial kinetic energy of the shell is equal to its potential energy at maximum height:


"\dfrac{mv^2}{2} = mgh"

where "v = 100m\/s, g = 9.8m\/s^2," "m" is the mass of the shell, and "h" is the maximum height. Expressing "h", obtain:


"h = \dfrac{v^2}{2g} = \dfrac{100^2}{2\cdot 9.8} \approx 510m"

Answer. 510 m.